OpenStudy (anonymous):

If AC=4x+5 and BD=x+23, what is the value of x? I

Mehek (mehek14):

do you have an attachment?

OpenStudy (anonymous):

Um no, but it is basically this but backwards.

OpenStudy (owlcoffee):

Well, if they are diagonals, they should be equal, It would be long to prove, but it is an "isoceles trapezoid" because it has two pairs of equal angles, wich would mean it has both sides equal (those that are diagonal to the sup.base and the inf. base). As I said, in all isoceles trapezoids, the diagonals are equal. So we will use that information: \[AC=BD\] Therefore: \[4x+5=x+23\] And it transformed into an univariable equation, wich we can solve by moving all the variables in one side of the equality and the constants to the other: \[4x-x=23-5\] And with the magic of aritmethics: \[4x-x=18\] And operating similar variables: \[3x=18\] And then, isolating "x": \[x= \frac{ 18 }{ 3 }\] And there you have it, all you have to do is see what 18/3 is equal to, and you'll be done.

OpenStudy (anonymous):

what do you mean see what 18/3 is equal to?